Show the derivation of the formula.


Cos (u+v)=cos u cos v - sin u sin v

1
Expert's answer
2022-03-02T03:18:48-0500

Derivation of the formula.

Cos (u+v)=cos u cos v - sin u sin v






(QP)2 = (cos U - Cos V)2 + (sin U - sin V)2

= cos2 u - 2 cos u cos v + cos2v + sin2u -2 sin u sin v +sin2v

= 2 - 2(cos u cos v +sin u sin v) ...................................................(eq1)

(QP)2 = 12 + 12 -2 *1 *1 * cos (u-v)....(cosine formula)

= 2 - 2 cos (u -v)............................................................................(eq2)


Equating eq1 and eq2 we get,


cos (u - v) = cos u cos v +sin u sin v


Let v = -v


cos(u−(−v))=cos u cos(−v) + sinu sin(−v)


(since cos (-a) = cos a and sin (-a) = - sin a)


hence

cos (u + v) = cos u cos v - sin u sin v [Answer]